OpenStudy (stickman1998):

Can someone show me how to do this, please? Verify the Pythagorean identity.

OpenStudy (stickman1998):

\[1+\cot^2 \theta = \csc^2 \theta\]

OpenStudy (mathstudent55):

Add the two terms on the left. Write the cotangent as a fraction, find a common denominator, and add them together. Then pay close attention to the numerator of the sum.

OpenStudy (stickman1998):

Umm, do you think you could simplify that at all?

OpenStudy (ethan01):

Here are all the steps in one post 1+cot2(θ)=csc2(θ) 1+cos2(θ)sin2(θ)=csc2(θ) sin2(θ)sin2(θ)+cos2(θ)sin2(θ)=csc2(θ) sin2(θ)+cos2(θ)sin2(θ)=csc2(θ) 1sin2(θ)=csc2(θ) csc2(θ)=csc2(θ) the two sides of the last equation are identical, so we have confirmed the identity

OpenStudy (mathstudent55):

\(1+\cot^2 \theta = \csc^2 \theta\) \(1 + \dfrac{\cos^2 \theta}{\sin^2 \theta} = \csc^2 \theta\) We need a common denominator to add the left side. \(\dfrac{\sin^2 \theta}{\sin^2 \theta}+ \dfrac{cos^2 \theta}{\sin^2 \theta} = \csc^2 \theta\) \(\dfrac{\sin^2 \theta + cos^2 \theta}{\sin^2 \theta} = \csc^2 \theta\) Use the identity: \(\sin^2 \theta + \cos^2 \theta = 1\) \(\dfrac{1}{\sin^2 \theta} = \csc^2 \theta\) \(\csc^2 \theta = \csc^2 \theta\)

OpenStudy (stickman1998):

Awesome! Thanks a bunch for explaining it

OpenStudy (mathstudent55):

You're welcome.